OpenStudy (anonymous):

This is the hardest question yet! During the basketball season, Diane took 134 shots and made about 56% of them. a. How many shots did Diane make? b. The team made a total of 498 shots. What percent of the team’s made shots did Diane make? 24 shots; 15.1% 753 shots; 27% 240 shots; 27% 75 shots; 15.1% Thanks Again every one +D

OpenStudy (anonymous):

here: http://openstudy.com/study#/updates/509aa028e4b02ec0829d2195

OpenStudy (anonymous):

wow.. dude just take your percentage into a decimal and multiply it to the main number a) 56% = 0.56 0.56(134)= N N=final answer for (A) b) N/498 = T T(100) = X X=final answer for (B)

OpenStudy (anonymous):

You know this question has been asked twice

OpenStudy (anonymous):

Diane shots= 0.56 * 134 = 75.04 = 75 as it is written about %age of diane shots with reference to team = 75*100/498= 15.1%